User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!



Power levels on last 2 channels of iread (s-band) keep fluctuating and can't stay near target for long. Jim thinks it's probably RFI so we can't do anything about it at the moment. Brett should be able to investigate when he gets in tomorrow. (Ellen)

Module: HOB+0120, start pos 19384.631 GB

  • 18:30 UT - Experiment started ok. (Ellen)


Dimino wouldn't run properly. Rebooted mark 5. It wouldn't let us ssh in but we could ping the server. Jim eventually got it started again. (Ellen)

Module: IPA-0022, start pos 0 GB

  • 18:30 UT - Experiment started ok. (Ellen)
  • 20:45 UT - All tsys values overflowing for about 15 minutes. There is no storm to interfere and this stretches across multiple scans so is not position dependent. Continuing to watch. (Ellen)
  • 22:00 UT - Tsys values are still all overflowing, Jim thinks there's nothing wrong with the data just the values it's displaying. He will investigate at 9am. (Ellen)
  • 22:30 UT - Every few scans the autocorrelation shows a jump in band 5 - noted in log at specific times. (Ellen)
  • 16:15 UT - Channel 4 of the autocorrelation looked like a “step function” and all the system temps had disappeared. Reconfigured the dbbc between the 16:15 and 16:20 scans. No scan should be affected, auto correl spectra back to normal. Also updated the wiki page on this. (AK)


Module: NTO-0007, start pos 0 GB

  • 18:30 UT - Experiment started ok. (Ellen)
You could leave a comment if you were logged in.
/home/www/auscope/opswiki/data/pages/handover/r4757.txt · Last modified: 2016/09/16 16:41 by Jesse Swan